# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v3

@@ 3.5 Hyperparameter independence
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent for $\tilde{g}=sBg^{R}A$, but not for $\tilde{g}=sBg^{R}_{LoRA-XS}A$
@@
See Appendix A.5. ∎
The hyperparameter independence of the equivalent gradient eliminates the need for manual gradient scaling. Updates to $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant. This can be understood by examining the relationship with the full FT gradient $g$. Since $g$ is naturally scaled for optimal weight updates, and our method approximates $g$ in a constrained subspace, the equivalent gradient inherits appropriate scaling automatically. This property is unique to our gradient approximation approach and does not hold for standard LoRA-XS.
@@ 3.6 LoRA-SB: Update approximation initialization is a silver bullet
The solutions discussed above independently address the gradient approximation and initialization problems, while also providing hyperparameter independence. Our proposed method, LoRA-SB, elegantly combines these solutions through a simple initialization strategy, derived from approximating the first step of full FT:
@@
$B_{init}\leftarrow U[1:r]$ (9)
$A_{init}\leftarrow V[1:r]$ (10)
$R_{init}\leftarrow\dfrac{1}{s}S[1:r,1:r]$ (11)
By the Eckart-Young theorem (Eckart & Young, 1936; Mirsky, 1960), this gives the optimal rank-$r$ approximation of the full FT update. where $U$, $S$, $V$ are obtained from truncated SVD of the averaged first update $\Delta W_{\text{avg}}$. This initialization leads to several key advantages.
@@
$g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}.$
Optimal Update Approximation. Our initialization guarantees that the first update optimally approximates the full FT weight updates:
$sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}.$
By the Eckart-Young theorem, this gives the optimal rank-$r$ approximation of the initial full FT update.
Hyperparameter Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in:
@@
This demonstrates that our initialization guarantees optimal gradient approximation at every step, without requiring any scaling factor.
Guaranteed Loss Reduction. Since $B$ is a tall orthonormal matrix and $A$ is a wide orthonormal matrix, they remain full rank throughout training. This ensures that $dL$ remains negative 4, guaranteeing stable optimization and convergence.
Another heuristic which might lead to a good initialization is setting the weights $B$ and $A$, such that they match the first update also approximately matches the direction of $\Delta W$.
@@
Thankfully, we don’t have to choose between the two. For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 8-11, results in the first update of LoRA-XS to best approximate the direction of the update of full FT (Theorem 6).
@@
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
@@
See Appendix A.6. ∎
LoRA-SB Advantages over LoRA. Many properties described above are not achievable with standard LoRA methods. Even if $B$ and $A$ are initialized as orthonormal in LoRA, subsequent updates do not preserve this property because $B$ and $A$ are trainable. This results in several challenges:
  • Potential instability of $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$, as they are not guaranteed to remain non-singular during training.
  • Inability to ensure consistent loss reduction due to potential rank deficiency—$B$ and $A$ may not remain full-rank throughout training.
  • Necessity to fine-tune the hyperparameter $\alpha$.
@@
Algorithm. We present a PyTorch-like implementation of our method in Algorithm 1. To optimize GPU memory usage during initialization, we hook into the backward pass of PyTorch and compute the gradients layerwise, immediately discarding the computed gradients (Lv et al., 2024; Wang et al., 2024a). This ensures $O(1)$ memory usage, independent of the number of layers, keeping memory consumption well within manageable limits and ensuring it does not exceed the memory requirements of subsequent LoRA-SB FT. For large batch sizes, memory usage can be further optimized through gradient accumulation and quantization.
Algorithm 1 LoRA-SB, PyTorch-like
  def initSB(model, D)
       # Estimate gradient with n samples
       $\mathbf{\Delta W_{avg}}\leftarrow$ est_grad(model, D, n)
       # Initialize B, R, A
       $\mathbf{B,R,A}\leftarrow$ trunc_SVD($\mathbf{\Delta W_{avg}}$)
       # Convert to LoRA-SB model
[…]
       return sb_model
  # Load pre-trained model
[…]
  # Initialize LoRA-SB with dataset D
[…]
  # Standard training, only R trainable
[…]
@@ 4 Experiments
We evaluate our method over $16$ different datasets on three widely-used NLP benchmarks, using models ranging from the 355 M-parameter RoBERTa-large model to the 9 B-parameter Gemma-2 model. Our setup spans both masked and autoregressive architectures, allowing us to comprehensively assess the effectiveness of our approach. We fine-tune RoBERTa-large (Liu et al., 2019), Llama-3.2 3B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Gemma-2 9B (Team et al., 2024), showcasing our method’s adaptability across a variety of tasks and model architectures.
We use PyTorch (Paszke et al., 2019) and the HuggingFace Transformers library (Wolf et al., 2020) for our implementations. We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds. To save memory, we initialize base models in torch.bfloat16 precision. Appendix D provides detailed information on the datasets used. We compute the update approximation using only $\mathbf{1/1000}$ of each dataset’s total number of samples. This ensures that the additional training time overhead is minimal and has a negligible effect on overall efficiency. We provide a benchmark of this overhead in Section 5. The samples are randomly selected from the training set in each run. Detailed hyperparameter settings for all tasks are given in Appendix C.
Baselines. We compare LoRA-SB against full FT, LoRA, LoRA-XS, rsLoRA, and PiSSA. rsLoRA introduces a rank-scaled stabilization factor ($\alpha/\sqrt{r}$) to enhance stability, while PiSSA updates only the principal components of the pre-trained weight $W$ and freezes the residuals.
4.1 Arithmetic Reasoning
@@
Table 1: Accuracy comparison of FT methods on Mistral-7B and Gemma-2 9B across the arithmetic reasoning benchmarks GSM8K and MATH, after training on MetaMathQA. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
Model Method Rank # Params Accuracy ($\uparrow$)
GSM8K MATH
Mistral-7B Full FT - $7.24$ B $63.87$ $17.65$
LoRA $32$ $83.88$ M $61.94$ $15.98$
rsLoRA $32$ $83.88$ M $62.15$ $16.24$
PiSSA $32$ $83.88$ M $62.43$ $16.52$
LoRA-XS $32$ $0.23$ M $54.28$ $13.36$
LoRA-XS $64$ $0.92$ M $57.08$ $15.62$
LoRA-XS $96$ $2.06$ M $58.53$ $16.42$
LoRA-SB $32$ $0.23$ M $58.91$ $15.28$
LoRA-SB $64$ $0.92$ M $60.73$ $16.28$
LoRA-SB $96$ $2.06$ M $\mathbf{63.38}$ $\mathbf{17.44}$
Gemma-2 9B Full FT - $9.24$ B $79.23$ $38.02$
LoRA $32$ $108.04$ M $76.19$ $36.56$
rsLoRA $32$ $108.04$ M $76.84$ $36.88$
PiSSA $32$ $108.04$ M $77.12$ $37.04$
LoRA-XS $32$ $0.30$ M $74.07$ $34.62$
LoRA-XS $64$ $1.20$ M $75.02$ $36.46$
LoRA-XS $96$ $2.71$ M $75.21$ $36.98$
LoRA-SB $32$ $0.30$ M $75.44$ $36.66$
LoRA-SB $64$ $1.20$ M $76.65$ $37.14$
LoRA-SB $96$ $2.71$ M $\mathbf{78.40}$ $\mathbf{37.70}$
Details. We fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 50K samples from the MetaMathQA (Yu et al., 2024) dataset and evaluate them on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks. We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
Results. We present results in Table 1. LoRA-SB significantly outperforms LoRA-XS across all settings. Notably, LoRA-SB outperforms LoRA-based methods ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively. We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2. Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS. Additionally, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss curve throughout and converges to a superior final value.
(a) Mistral-7B
(b) Gemma-2 9B
Figure 2: Training loss curves for Mistral-7B and Gemma-2 9B on MetaMathQA, comparing LoRA-SB and LoRA-XS.
4.2 Commonsense Reasoning
Details. We fine-tune Llama-3.2 3B (Dubey et al., 2024) on CommonSense170K, a dataset with eight commonsense reasoning tasks (Hu et al., 2023). We evaluate the model’s performance on each dataset individually, which include BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019). LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
Results. We present the results in Table 2. LoRA-SB consistently outperforms LoRA-XS across all settings. In addition, LoRA-SB ($r=96$) outperforms LoRA-based methods ($r=32$) with 27x fewer trainable parameters.
@@ 4.2 Commonsense Reasoning
Table 2: Accuracy comparison of FT methods on Llama-3.2 3B across eight commonsense reasoning datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
Model Method Rank # Params Accuracy ($\uparrow$)
@@
rsLoRA $32$ $48.63$ M $69.81$ $85.05$ $78.92$ $90.45$ $82.02$ $86.71$ $74.18$ $81.72$ $81.11$
PiSSA $32$ $48.63$ M $70.12$ $\mathbf{85.42}$ $79.44$ $90.88$ $82.68$ $87.23$ $74.61$ $81.79$ $81.52$
@@
LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $\mathbf{74.83}$ $80.12$ $81.29$
LoRA-SB $96$ $1.81$ M $\mathbf{70.34}$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $74.74$ $81.20$ $\mathbf{81.92}$
4.3 Natural Language Understanding
Details. We fine-tune RoBERTa-large (Liu et al., 2019) on GLUE, a popular language understanding classification benchmark that contains several datasets. The datasets we evaluate on are: CoLA (Warstadt et al., 2019), RTE, MRPC (Dolan & Brockett, 2005), SST-2 (Socher et al., 2013), QNLI (Rajpurkar et al., 2018), and STS-B (Cer et al., 2017). LoRA modules are applied only to the self-attention layers, with ranks $r=\{8,16,24\}$.
Results. The results are shown in Table 3. LoRA-SB consistently outperforms LoRA-XS across all configurations. Additionally, LoRA-SB ($r=24$) outperforms LoRA-based methods ($r=8$) with 39x lesser trainable parameters.
Table 3: Comparison of FT methods on RoBERTa-large across the GLUE benchmark datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold. We use Pearson correlation for STS-B, Matthew’s correlation for CoLA, and accuracy for others. Higher is better for each metric.
Model Method Rank # Params CoLA RTE MRPC STS-B QNLI SST-2 All
@@ 4.3 Natural Language Understanding
RoBERTa-large Full FT - $355.36$ M $68.44$ $83.42$ $90.21$ $91.76$ $93.92$ $96.21$ $87.33$
LoRA $8$ $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
rsLoRA $8$ $2162.69$ K $67.87$ $82.84$ $89.97$ $91.30$ $93.29$ $95.87$ $86.85$
PiSSA $8$ $2162.69$ K $68.22$ $\mathbf{83.14}$ $90.10$ $91.59$ $93.55$ $96.03$ $87.10$
LoRA-XS $8$ $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
@@
LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $83.03$ $\mathbf{90.12}$ $\mathbf{91.65}$ $93.75$ $\mathbf{96.11}$ $\mathbf{87.16}$

### v4

@@ 2.5 Scaling Factor independence
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent for $\tilde{g}=sBg^{R}A$, but not for $\tilde{g}=sBg^{R}_{LoRA-XS}A$.
@@
See Appendix B.5. ∎
The scaling factor independence of the equivalent gradient eliminates the need for manual gradient scaling. Updates to $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant. This can be understood by examining the relationship with the full FT gradient $g$. Since $g$ is naturally scaled for optimal weight updates, and our method approximates $g$ in a constrained subspace, the equivalent gradient inherits appropriate scaling automatically. This property is unique to our gradient approximation approach and does not hold for standard LoRA-XS.
@@ 2.6 LoRA-SB: Update approximation initialization is a silver bullet
The solutions discussed independently address the gradient approximation and initialization problems, while also providing scaling factor independence. LoRA-SB, elegantly combines these solutions through a simple initialization strategy, derived from approximating the first full FT step:
@@
$B_{init}\leftarrow U[1:r],A_{init}\leftarrow V[1:r],R_{init}\leftarrow\dfrac{1}{s}S[1:r,1:r]$ (8)
By the Eckart-Young theorem (13, 32), this gives the optimal rank-$r$ approximation of the full FT update. where $U$, $S$, $V$ are obtained from truncated SVD of the averaged first update $\Delta W_{\text{avg}}$. This initialization leads to several key advantages.
@@
$g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}$
Optimal Update Approximation. Our initialization guarantees that the first update optimally approximates the full FT weight updates: $sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}.$ By the Eckart-Young theorem, this is the optimal rank-$r$ approximation of the initial full FT update.
Scaling Factor Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in guaranteed optimal gradient approximation at every step, without requiring any scaling factor:
@@
Guaranteed Loss Reduction. Since $B$ is a tall orthonormal and $A$ a wide orthonormal matrix, they remain full rank throughout training. This ensures that $dL$ remains negative (Theorem 4), guaranteeing stable optimization and convergence.
@@
Another heuristic which might lead to a good initialization is setting $B$ and $A$, such that the first update also approximately matches the $\Delta W$ direction (Equation 10). Thankfully, we don’t have to choose between the two. For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 7-8, results in the first update of LoRA-XS to best approximate the direction of the full FT update (Theorem 6).
@@
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then the update given by LoRA-SB, $\Delta(B_{init}R_{init}A_{init})$ , is the best low-rank approximation of full fine-tuning update, $\Delta W$.
@@
See Appendix B.6. ∎
While Theorem 6 is stated for SGD, the result extends to other SGD-based optimizers such as AdamW. In practice, we use AdamW and approximate the first update by taking the sign of the averaged gradients, consistent with AdamW’s first-step behavior. This produces an initialization whose SVD still yields the optimal rank-$r$ approximation of the simulated full FT update.
Initialization Memory. To optimize GPU memory during initialization, we hook into the backward pass and compute the gradients layerwise, immediately discarding the computed gradients (29, 45). This ensures $O(1)$ memory usage, independent of the number of layers, keeping GPU memory well within limits. This guarantees that the memory required for LoRA-SB initialization never exceeds the memory needed for subsequent LoRA-SB fine-tuning, and that the peak memory usage of the entire LoRA-SB algorithm never exceeds that of standard LoRA and other baselines.
LoRA-SB Advantages over LoRA. Many properties described above are not achievable with standard LoRA methods. Even if $B$ and $A$ are initialized as orthonormal in LoRA, subsequent updates do not preserve this property because $B$ and $A$ are trainable. This results in several challenges in using LoRA (even with optimal gradient approximation) compared to LoRA-SB:
  • Potential instability of $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$, not guaranteed to remain non-singular throughout.
  • Inability to ensure consistent loss reduction due to potential rank deficiency, $B$ and $A$ may not remain full-rank throughout training.
  • Necessity to fine-tune the scaling factor hyperparameter $\alpha$.
@@ 3 Experiments
We evaluate over $16$ different datasets on $3$ widely-used benchmarks, using models ranging from the 355 M RoBERTa-large model to the 9 B Gemma-2 model. Our setup spans both masked and autoregressive architectures, allowing us to comprehensively assess the effectiveness of LoRA-SB. Specifically, we fine-tune RoBERTa-large (27), Llama-3.2 3B (12), Mistral-7B (19), and Gemma-2 9B (43). We compute the update approximation using only $\mathbf{1/1000}$ ($0.1\%$) of each dataset’s total size. This ensures that the training time overhead is minimal and has a negligible effect on efficiency. Detailed hyperparameter and dataset details are given in Appendix H and I, respectively.
Baselines. We compare LoRA-SB against full FT, LoRA (17), LoRA-XS (2), and several popular variants of LoRA - rsLoRA (20), PiSSA (30), DoRA (26), and LoRA-Pro (46).
@@
Table 1: Comparison of FT methods on Mistral-7B and Gemma-2 9B across arithmetic benchmarks. # Params denotes the number of trainable parameters. Best results among PEFT methods are in bold.
Method Rank Mistral-7B Gemma-2 9B
# Params GSM8K ($\uparrow$) MATH ($\uparrow$) # Params GSM8K ($\uparrow$) MATH ($\uparrow$)
Full FT - $7.24$ B $63.87$ $17.65$ $9.24$ B $79.23$ $38.02$
LoRA 32 $83.88$ M $61.94$ $15.98$ $108.04$ M $76.19$ $36.56$
rsLoRA 32 $83.88$ M $62.15$ $16.24$ $108.04$ M $76.84$ $36.88$
PiSSA 32 $83.88$ M $62.43$ $16.52$ $108.04$ M $77.12$ $37.04$
DoRA 32 $85.26$ M $62.65$ $16.64$ $109.88$ M $77.58$ $37.04$
LoRA-Pro 32 $83.88$ M $63.07$ $17.32$ $108.04$ M $78.26$ $37.53$
LoRA-XS 32 $0.23$ M $54.28$ $13.36$ $0.30$ M $74.07$ $34.62$
LoRA-XS 64 $0.92$ M $57.08$ $15.62$ $1.20$ M $75.02$ $36.46$
LoRA-XS 96 $2.06$ M $58.53$ $16.42$ $2.71$ M $75.21$ $36.98$
LoRA-SB 32 $0.23$ M $58.91$ $15.28$ $0.30$ M $75.44$ $36.66$
LoRA-SB 64 $0.92$ M $60.73$ $16.28$ $1.20$ M $76.65$ $37.14$
LoRA-SB 96 $2.06$ M $\mathbf{63.38}$ $\mathbf{17.44}$ $2.71$ M $\mathbf{78.40}$ $\mathbf{37.70}$
@@
Table 2: Comparison of FT methods on Llama-3.2 3B across eight commonsense reasoning datasets. # Params denotes the number of trainable parameters. Best results among PEFT methods are in bold.
Method Rank # Params Accuracy ($\uparrow$)
@@
rsLoRA $32$ $48.63$ M $69.81$ $85.63$ $78.92$ $90.45$ $82.02$ $86.71$ $74.18$ $81.72$ $81.11$
PiSSA $32$ $48.63$ M $70.12$ $85.42$ $79.44$ $90.88$ $82.68$ $87.23$ $74.61$ $81.79$ $81.52$
DoRA $32$ $49.40$ M $70.43$ $85.63$ $79.68$ $90.76$ $82.90$ $87.61$ $74.87$ $82.04$ $81.74$
LoRA-Pro $32$ $48.63$ M $\mathbf{71.28}$ $\mathbf{85.81}$ $79.35$ $90.90$ $83.42$ $87.24$ $\mathbf{75.32}$ $81.74$ $81.88$
@@
LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $74.83$ $80.12$ $81.29$
LoRA-SB $96$ $1.81$ M $70.34$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $74.74$ $81.20$ $\mathbf{81.92}$
Table 3: Comparison of FT methods on RoBERTa-large across GLUE datasets. # Params denotes the number of trainable parameters. Best results among PEFT methods are in bold. We use Pearson correlation for STS-B, Matthew’s correlation for CoLA, and accuracy for others.
Method Rank # Params CoLA RTE MRPC STS-B QNLI SST-2 All
@@
Full FT - $355.36$ M $68.44$ $83.42$ $90.21$ $91.76$ $93.92$ $96.21$ $87.33$
LoRA 88 $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
rsLoRA 88 $2162.69$ K $67.87$ $82.84$ $89.97$ $91.30$ $93.29$ $95.87$ $86.85$
PiSSA 88 $2162.69$ K $68.22$ $83.14$ $90.10$ $91.59$ $93.55$ $96.03$ $87.10$
DoRA 88 $2260.99$ K $68.05$ $83.04$ $89.93$ $91.34$ $93.11$ $95.82$ $86.88$
LoRA-Pro 88 $2162.69$ K $67.98$ $\mathbf{83.40}$ $\mathbf{90.49}$ $91.38$ $93.37$ $95.98$ $87.10$
LoRA-XS 88 $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
@@
LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $83.03$ $90.12$ $\mathbf{91.65}$ $93.75$ $\mathbf{96.11}$ $\mathbf{87.16}$
3.1 Arithmetic Reasoning
We fine-tune Mistral-7B (19) and Gemma-2 9B (43) on 50K samples from MetaMathQA (50) and evaluate on GSM8K (8) and MATH (16). We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$. We present results in Table 1. LoRA-SB significantly outperforms LoRA-XS across all settings. LoRA-SB outperforms LoRA-based methods ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively. We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2. Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS. Further, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss throughout and converges to a superior final value.
(a) Mistral-7B
(b) Gemma-2 9B
Figure 2: Training loss curves for Mistral-7B and Gemma-2 9B, comparing LoRA-SB and LoRA-XS.
3.2 Commonsense Reasoning
We fine-tune Llama-3.2 3B (12) on CommonSense170K, a dataset with eight commonsense reasoning tasks (18). LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$. We present the results in Table 2. LoRA-SB consistently outperforms LoRA-XS across all settings. In addition, LoRA-SB ($r=96$) outperforms LoRA-based methods ($r=32$) with 27x fewer trainable parameters.
3.3 Natural Language Understanding
We fine-tune RoBERTa-large (27) on GLUE, a popular language understanding benchmark. LoRA modules are applied only to the self-attention layers, with ranks $r=\{8,16,24\}$. Results are shown in Table 3. LoRA-SB consistently outperforms LoRA-XS across all settings. Additionally, LoRA-SB ($r=24$) outperforms LoRA-based methods ($r=8$) with 39x lesser trainable parameters.